# Self-Speculative Masked Diffusions

### v1

@@ Abstract
We present self-speculative masked diffusions, a new class of masked diffusion generative models for discrete data that require significantly fewer function evaluations to generate samples. Standard masked diffusion models predict factorized logits over currently masked positions. A number of masked positions are then sampled, however, the factorization approximation means that sampling too many positions in one go leads to poor sample quality. As a result, many simulation steps and therefore neural network function evaluations are required to generate high-quality data. We reduce the computational burden by generating non-factorized predictions over masked positions. This is achieved by modifying the final transformer attention mask from non-causal to causal, enabling draft token generation and parallel validation via a novel, model-integrated speculative sampling mechanism. This results in a non-factorized predictive distribution over masked positions in a single forward pass. We apply our method to GPT2 scale text modelling and protein sequences generation, finding that we can achieve a $\sim 2\times$ reduction in the required number of network forward passes relative to standard masked diffusion models.
@@ 1 Introduction
Generative models for discrete data are at the core of a wide range of modern deep learning systems from chatbots (Gemini Team,, 2023; OpenAI Team,, 2023; Anthropic,, 2024) to biological foundation models (Hayes et al.,, 2025; Wang et al.,, 2025; Madani et al.,, 2023). These models generate data through an iterative process, each step revealing only a small portion of the final tokens. Autoregressive (AR) models (Sutskever et al.,, 2014; Brown et al.,, 2020) reveal a single token in each generation step, doing so in a left-to-right ordering. Alternatively, masked diffusions models (MDMs) and any-order AR models (Hoogeboom et al.,, 2022; Shih et al.,, 2022; Sahoo et al.,, 2024; Shi et al.,, 2024; Ou et al.,, 2025) reveal multiple tokens at each step and can operate using any given generation ordering. The ordering flexibility is beneficial for applications without an inherent left-to-right structure such as protein sequences (Alamdari et al.,, 2023; Gruver et al.,, 2024; Wang et al.,, 2024).
During an update, standard MDMs use a neural network to output a factorized predictive distribution over currently masked positions. Revealing more than one token from this distribution incurs approximation error, as data distributions typically do not factorize that way; see e.g. (Zheng et al.,, 2025). This imposes an upper limit on how many tokens can be revealed per step without loss of quality, meaning a large number of network forward passes are required to generate a full datapoint.
Our objective is to reveal multiple masked tokens concurrently using a non-factorized predictive distribution. Inspired by self-speculative sampling (Zhang et al.,, 2024; Elhoushi et al.,, 2024; Liu et al.,, 2024), we propose generating a draft sequence using a network subset, then validating it in parallel with the full transformer. By employing speculative sampling (Leviathan et al.,, 2023; Chen et al.,, 2023), we ensure the accepted token sequence is distributed according to a non-factorized target distribution defined by the full capacity network.
Using self-speculative sampling for MDMs provides a non-factorized predictive distribution that can be sampled efficiently, but requires solving two key challenges. Firstly, speculative verification requires a causal transformer but standard MDMs use a non-causal transformer. We solve this with a novel hybrid non-causal and permutation informed causal architecture (Pannatier et al.,, 2024), designing the information flow so that the causal target distribution is a strict improvement over the non-causal draft distribution. Second, in contrast to standard speculative sampling (Leviathan et al.,, 2023), the non-causal layers create a target distribution which depends on the speculative sampling accept/reject sequence. We theoretically characterize this dependence by deriving a log-likelihood lower bound for this model class.
@@ 2.1 Masked Diffusions as Any-Order Autoregressive Models
We assume our data $\bm{x}=(\bm{x}^{1},...,\bm{x}^{D})$ has $D$ dimensions and $S$ possible categories, i.e. $\bm{x}\in\{1,\dots,S\}^{D}$. Given the values of $\bm{x}$ on some subset of these dimensions, masking style generative models train a neural network to predict the remaining unobserved dimensions. This can be expressed mathematically in two ways. The first, MDMs, define a continuous-time discrete-valued Markov chain $(\bm{x}_{t}^{1:D})_{t\in[0,1]}$ on the state-space $\{1,\dots,S,S+1\}^{D}$, where an additional mask token $M$, $M:=S+1$, is introduced; see e.g. (Ou et al.,, 2025; Sahoo et al.,, 2024; Shi et al.,, 2024) for details. This Markov chain, initialized at $\bm{x}^{d}_{0}=\bm{x}^{d}$, is such that
$p(\bm{x}_{t}^{d}|\bm{x}^{d})=\alpha_{t}\updelta\{\bm{x}_{t}^{d}=\bm{x}^{d}\}+(1-\alpha_{t})\updelta\{\bm{x}_{t}^{d}=M\},$ (1)
for a non-increasing noising schedule $\alpha_{t}\in[0,1]$ such that $\alpha_{0}=1,~\alpha_{1}=0$. The time-reversal of this noising process defines a generative model consisting of $D$ random events spaced between $t=0$ and $t=1$, with each event revealing a single masked token. Individually simulating these $D$ events is computationally prohibitive when $D$ is large. Hence, many inference techniques instead reveal multiple masked tokens simultaneously using a factorized predictive distribution over the masked dimensions given the current partially masked context $\bm{x}_{t}^{1:D}$. This is parameterized by a neural network with parameters $\theta$ as,
@@
As shown by Austin et al., (2021); Hoogeboom et al., (2022); Ou et al., (2025), MDMs are closely related to any-order AR models which directly consider the unmasking events, doing away with the time variable $t$. Let $\sigma$ be a permutation of the numbers $\{1,\dots,D\}$, with $\sigma(i:j)=(\sigma(i),\sigma(i+1),...,\sigma(j))$ representing the $i$-th to the $j$-th element of this permutation. This permutation is usually picked uniformly at random. Given some subset of the dimensions of $\bm{x}$, any-order AR models predict the remaining dimensions by parameterizing the predictive distribution $p_{\theta}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$. AR diffusion models (Hoogeboom et al.,, 2022) use a factorized parameterization of this distribution similar to MDMs
@@
An alternative approach is to parameterize the predictive distribution autoregressively, using a network architecture that can operate over a given input ordering (Yang et al.,, 2019; Pannatier et al.,, 2024), i.e.
@@ 2.2 Speculative Sampling
Speculative sampling is a method developed to accelerate inference for large language models (LLMs) (Chen et al.,, 2023; Leviathan et al.,, 2023). The total cost of LLM inference is memory-bound, i.e., dominated by the cost of moving the model weights from high-bandwidth memory to the accelerator cache. Since cost is dominated by memory operations, applying the transformer to a sequence has similar latency to a single token, despite increased arithmetic computations in the former case. Speculative sampling exploits this by creating a draft sequence using a smaller model, which is then efficiently verified in parallel with a single forward pass of the larger target LLM. Verification results in the acceptance of a subset of the draft tokens, such that the accepted sequence is still distributed according to the probability distribution defined by the target model.
Formally, let the draft model be $p(\bm{x}^{d}|\bm{x}^{1:d-1})$ and the target model be $q(\bm{x}^{d}|\bm{x}^{1:d-1})$. We aim to draw samples from $q$ by first sampling a sequence of tokens from $p$ and accepting some number of them. We let the context of already generated tokens be $\bm{x}^{1:i}$ and let $W\geq 1$ be some window size. To sample new tokens, we first draw the draft sequence, $\smash{\hat{\bm{x}}^{i+1:i+W}\sim\prod_{d=i+1}^{i+W}p(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})}$. We then, in parallel, obtain the probabilities that the target model assigns to the draft tokens $\smash{\hat{\bm{x}}^{i+1:i+W}}$, $\{q(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})\}_{d=i+1}^{i+W}$ . Starting from the $(i+1)$-th token, we accept each token with probability $\text{min}\left(1,q(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})/p(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})\right)$. Let the first position we reject be $j$. We then resample $\hat{\bm{x}}^{j}$ from the adjusted distribution $r(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})\propto\text{max}\left(0,q(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})-p(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})\right)$. One can prove that the resulting sequence $\hat{\bm{x}}^{i+1:j}$ is distributed according to the target distribution $q(\hat{\bm{x}}^{i+1:j}|\bm{x}^{1:i})$ (Leviathan et al.,, 2023).
@@ 3.1 Architecture
Figure 1: Hybrid non-causal/causal transformer during training on the sentence “Speculation is like hazarding a guess” which is partially corrupted in the ordering given by $\sigma(1:6)=[6,5,2,4,3,1]$. We always mask the final elements in the sequence, here the final 33 elements in the sequence, $[4,3,1]$. At inference time, the causal transformer uses draft tokens rather than real tokens.
We need to parameterize both a non-causal factorized distribution $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$ and a causal AR distribution $\overset{\rightarrow}{p}_{\phi}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$ for any permutation $\sigma$. We propose a hybrid architecture that starts with a set of non-causal blocks parameterizing $\smash{\overset{\leftrightarrow}{p}_{\theta}}$ similar to standard MDMs. We then use a set of causal blocks operating on the hidden states from the previous layers to parameterize $\smash{\overset{\rightarrow}{p}_{\phi}}$. These causal blocks follow the $\sigma$-GPT architecture (Pannatier et al.,, 2024), which models any ordering $\sigma$ by operating on the permuted sequence with additional positional encodings. Our full architecture is given in Figure 1, we now detail each component.
@@
Causal Blocks. We parameterize the causal distribution via the $\sigma$-GPT architecture: left-to-right attention masked causal blocks applied to the permuted sequence. Unlike non-causal blocks, the track corresponding to position $\sigma(j)$ predicts the token for the next position in the sequence, $\sigma(j+1)$. This is standard for causally masked transformers; since the full unmasked sequence is input, predicting the token $\sigma(j)$ from track $\sigma(j)$ would be trivial. Because we operate on any ordering $\sigma$, each track needs information on both its own position, $\sigma(j)$, and the next position $\sigma(j+1)$, provided via positional encodings (Pannatier et al.,, 2024). For example, in Figure 1, the $\sigma(5)$ track uses positional encodings for $\sigma(5)=3$ and $\sigma(6)=1$ to predict the token “Speculation” in position $\sigma(6)=1$
In our architecture, we also input the non-causal hidden states corresponding to the current and next position, following the same reasoning as for the positional encodings. For the first position, $\sigma(1)$, we set the causal distribution equal to the non-causal distribution as no extra information is available.
Output. At the causal output, for the prediction of token in position $\sigma(j+1)$, we use a residual connection adding the non-causal hidden state for position $\sigma(j+1)$ e.g. h1 for the token “Speculation” in Figure 1. This ensures the causal target strictly improves over the non-causal distribution. It also aligns the draft and target distributions, increasing the speculative acceptance rate. The causal output from our hybrid architecture depends on both the parameters of the non-causal blocks ($\theta$) and the causal blocks ($\phi$). We write this as $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\bm{x}^{\sigma(i+1:d-1)}))$ specifying that tokens $\bm{x}^{\sigma(1:i)}$ were processed by $\theta$ in the non-causal blocks, whilst future tokens $\phi(\bm{x}^{\sigma(i+1:d-1)})$ were input only into the causal blocks.
@@ 3.2 Training Objective
We note that in Equation 8, there are multiple causal distributions that approximate the same true conditional $p_{\rm{data}}(\bm{x}^{\sigma(d)}|\bm{x}^{\sigma(1:d-1)})$ for different values of $i$ which specifies the computation path taken to parameterize $\overset{\rightarrow}{p}_{\theta,\phi}$. We discuss this further in Section 3.3. We train $\theta$ and $\phi$ by jointly by maximizing the standard cross entropy loss with a weighting $\frac{D}{D-i}$ that normalizes by the number of masked positions DiD-i (Hoogeboom et al.,, 2022)
@@
where the expectation is taken over $p_{\rm{data}}(\bm{x})p(\sigma)p(i)$ with $\sigma$ sampled uniformly and $p(i)$ following a noising schedule similar to MDMs. The first part of the loss is mathematically equivalent to the MDM loss (Shi et al.,, 2024; Sahoo et al.,, 2024), as shown by Ou et al., (2025). The second part of the loss can be seen as a standard AR cross-entropy loss computed over the masked tokens under a random order (Uria et al.,, 2014). We highlight that evaluating both $\overset{\leftrightarrow}{p}_{\theta}$ and $\overset{\rightarrow}{p}_{\theta,\phi}$ for the objective takes only one single forward pass of the hybrid network. Training curves (Section 5.1) show that the causal distribution $\overset{\rightarrow}{p}_{\theta,\phi}$ achieves a lower loss than the non-causal draft distribution $\overset{\leftrightarrow}{p}_{\theta}$, as it can model missing tokens with a non-factorized distribution.

### v2

@@ Abstract
We present self-speculative masked diffusions, a new class of masked diffusion generative models for discrete data that require significantly fewer function evaluations to generate samples. Standard masked diffusion models predict factorized logits over currently masked positions. A number of masked positions are then sampled, however, the factorization approximation means that sampling too many positions in one go leads to poor sample quality. As a result, many simulation steps and therefore neural network function evaluations are required to generate high-quality data. We reduce the computational burden by generating non-factorized predictions over masked positions. This is achieved by modifying the final transformer attention mask from non-causal to causal, enabling draft token generation and parallel validation via a novel, model-integrated speculative sampling mechanism. This results in a non-factorized predictive distribution over masked positions in a single forward pass. We apply our method to GPT2 scale text modelling and protein sequence generation, finding that we can achieve a $\sim 2\times$ reduction in the required number of network forward passes relative to standard masked diffusion models.
@@ 1 Introduction
Generative models for discrete data are at the core of a wide range of modern deep learning systems from chatbots (Gemini Team, 2023; OpenAI Team, 2023; Anthropic, 2024) to biological foundation models (Hayes et al., 2025; Wang et al., 2025; Madani et al., 2023). These models generate data through an iterative process, each step revealing only a small portion of the final tokens. Autoregressive (AR) models (Sutskever et al., 2014; Brown et al., 2020) reveal a single token in each generation step, doing so in a left-to-right ordering. Alternatively, masked diffusion models (MDMs) and any-order AR models (Hoogeboom et al., 2022; Shih et al., 2022; Sahoo et al., 2024; Shi et al., 2024; Ou et al., 2025) reveal multiple tokens at each step and can operate using any given generation ordering. The ordering flexibility is beneficial for applications without an inherent left-to-right structure such as protein sequences (Alamdari et al., 2023; Gruver et al., 2024; Wang et al., 2024).
During an update, standard MDMs use a neural network to output a factorized predictive distribution over currently masked positions. Revealing more than one token from this distribution incurs approximation error, as data distributions typically do not factorize that way; see e.g. (Zheng et al., 2025). This imposes an upper limit on how many tokens can be revealed per step without loss of quality, meaning a large number of network forward passes are required to generate a full datapoint.
Our objective is to reveal multiple masked tokens concurrently using a non-factorized predictive distribution. Inspired by self-speculative sampling (Zhang et al., 2024; Elhoushi et al., 2024; Liu et al., 2024), we propose generating a draft sequence using a network subset, then validating it in parallel with the full transformer. By employing speculative sampling (Leviathan et al., 2023; Chen et al., 2023), we ensure the accepted token sequence is distributed according to a non-factorized target distribution defined by the full capacity network.
Using self-speculative sampling for MDMs provides a non-factorized predictive distribution that can be sampled efficiently, but requires solving two key challenges. Firstly, speculative verification requires a causal transformer but standard MDMs use a non-causal transformer. We solve this with a novel hybrid non-causal and permutation informed causal architecture (Pannatier et al., 2024), building in residual connections so that the causal target distribution learns to improve over the non-causal draft distribution. Second, in contrast to standard speculative sampling (Leviathan et al., 2023), the non-causal layers create a target distribution which depends on the speculative sampling accept/reject sequence. We theoretically characterize this dependence by deriving a log-likelihood lower bound for this model class.
@@ 2.1 Masked Diffusions as Any-Order Autoregressive Models
We assume our data $\bm{x}=(\bm{x}^{1},...,\bm{x}^{D})$ has $D$ dimensions and $S$ possible categories, i.e. $\bm{x}\in\{1,\dots,S\}^{D}$. Given the values of $\bm{x}$ on some subset of these dimensions, masking style generative models train a neural network to predict the remaining unobserved dimensions. This can be expressed mathematically in two ways. The first, MDMs, define a continuous-time discrete-valued Markov chain $(\bm{x}_{t}^{1:D})_{t\in[0,1]}$ on the state-space $\{1,\dots,S,S+1\}^{D}$, where an additional mask token $M$, $M:=S+1$, is introduced; see e.g. (Ou et al., 2025; Sahoo et al., 2024; Shi et al., 2024) for details. This Markov chain, initialized at $\bm{x}^{d}_{0}=\bm{x}^{d}$, is such that
$p(\bm{x}_{t}^{d}|\bm{x}^{d})=(1-\alpha_{t})\updelta\{\bm{x}_{t}^{d}=\bm{x}^{d}\}+\alpha_{t}\updelta\{\bm{x}_{t}^{d}=M\},$ (1)
for a non-decreasing noising schedule $\alpha_{t}\in[0,1]$ such that $\alpha_{0}=0,~\alpha_{1}=1$. The time-reversal of this noising process defines a generative model consisting of $D$ random events spaced between $t=0$ and $t=1$, with each event revealing a single masked token. Individually simulating these $D$ events is computationally prohibitive when $D$ is large. Hence, many inference techniques instead reveal multiple masked tokens simultaneously using a factorized predictive distribution over the masked dimensions given the current partially masked context $\bm{x}_{t}^{1:D}$. This is parameterized by a neural network with parameters $\theta$ as,
@@
As shown by Austin et al. (2021); Hoogeboom et al. (2022); Ou et al. (2025), MDMs are closely related to any-order AR models which directly consider the unmasking events, doing away with the time variable $t$. Let $\sigma$ be a permutation of the numbers $\{1,\dots,D\}$, with $\sigma(i:j)=(\sigma(i),\sigma(i+1),...,\sigma(j))$ representing the $i$-th to the $j$-th element of this permutation. This permutation is usually picked uniformly at random. Given some subset of the dimensions of $\bm{x}$, any-order AR models predict the remaining dimensions by parameterizing the predictive distribution $p_{\theta}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$. AR diffusion models (Hoogeboom et al., 2022) use a factorized parameterization of this distribution similar to MDMs
@@
An alternative approach is to parameterize the predictive distribution autoregressively, using a network architecture that can operate over a given input ordering (Yang et al., 2019; Pannatier et al., 2024), i.e.
@@ 2.2 Speculative Sampling
Speculative sampling is a method developed to accelerate inference for large language models (LLMs) (Chen et al., 2023; Leviathan et al., 2023). The total cost of LLM inference is memory-bound, i.e., dominated by the cost of moving the model weights from high-bandwidth memory to the accelerator cache. Since cost is dominated by memory operations, applying the transformer to a sequence has similar latency to a single token, despite increased arithmetic computations in the former case. Speculative sampling exploits this by creating a draft sequence using a smaller model, which is then efficiently verified in parallel with a single forward pass of the larger target LLM. Verification results in the acceptance of a subset of the draft tokens, such that the accepted sequence is still distributed according to the probability distribution defined by the target model.
Formally, let the draft model be $p(\bm{x}^{d}|\bm{x}^{1:d-1})$ and the target model be $q(\bm{x}^{d}|\bm{x}^{1:d-1})$. We aim to draw samples from $q$ by first sampling a sequence of tokens from $p$ and accepting some number of them. We let the context of already generated tokens be $\bm{x}^{1:i}$ and let $W\geq 1$ be some window size. To sample new tokens, we first draw the draft sequence, $\smash{\hat{\bm{x}}^{i+1:i+W}\sim\prod_{d=i+1}^{i+W}p(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})}$. We then, in parallel, obtain the probabilities that the target model assigns to the draft tokens $\smash{\hat{\bm{x}}^{i+1:i+W}}$, $\{q(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})\}_{d=i+1}^{i+W}$ . Starting from the $(i+1)$-th token, we accept each token with probability $\text{min}\left(1,q(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})/p(\hat{\bm{x}}^{d}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:d-1})\right)$. Let the first position we reject be $j$. We then resample $\hat{\bm{x}}^{j}$ from the adjusted distribution $r(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})\propto\text{max}\left(0,q(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})-p(\hat{\bm{x}}^{j}|\bm{x}^{1:i},\hat{\bm{x}}^{i+1:j-1})\right)$. One can prove that the resulting sequence $\hat{\bm{x}}^{i+1:j}$ is distributed according to the target distribution $q(\hat{\bm{x}}^{i+1:j}|\bm{x}^{1:i})$ (Leviathan et al., 2023).
@@ 3.1 Architecture
Figure 1: Hybrid non-causal/causal transformer during training on the sentence “Speculation is like hazarding a guess” which is partially corrupted in the ordering given by $\sigma(1:6)=[6,5,2,4,3,1]$. We always mask the final elements in the sequence, here the final 33 elements in the sequence, $[4,3,1]$. At inference time, the causal transformer uses draft tokens rather than real tokens. Empty circles represent concatenation whilst circles with a plus represent addition.
We need to parameterize both a non-causal factorized distribution $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$ and a causal AR distribution $\overset{\rightarrow}{p}_{\phi}(\bm{x}^{\sigma(i+1:D)}|\bm{x}^{\sigma(1:i)})$ for any permutation $\sigma$. We propose a hybrid architecture that starts with a set of non-causal blocks parameterizing $\smash{\overset{\leftrightarrow}{p}_{\theta}}$ similar to standard MDMs. We then use a set of causal blocks operating on the hidden states from the previous layers to parameterize $\smash{\overset{\rightarrow}{p}_{\phi}}$. These causal blocks follow the $\sigma$-GPT architecture (Pannatier et al., 2024), which models any ordering $\sigma$ by operating on the permuted sequence with additional positional encodings. Our full architecture is given in Figure 1, we now detail each component.
@@
Causal Blocks. We parameterize the causal distribution via the $\sigma$-GPT architecture: left-to-right attention masked causal blocks applied to the permuted sequence. Unlike non-causal blocks, the track corresponding to position $\sigma(j)$ predicts the token for the next position in the sequence, $\sigma(j+1)$. This is standard for causally masked transformers; since the full unmasked sequence is input, predicting the token $\sigma(j)$ from track $\sigma(j)$ would be trivial. Because we operate on any ordering $\sigma$, each track needs information on both its own position, $\sigma(j)$, and the next position $\sigma(j+1)$, provided via positional encodings (Pannatier et al., 2024). For example, in Figure 1, the $\sigma(5)$ track uses positional encodings for $\sigma(5)=3$ and $\sigma(6)=1$ to predict the token “Speculation” in position $\sigma(6)=1$
In our architecture, we also input the non-causal hidden states corresponding to the current and next position into the causal blocks, following the same reasoning as for the positional encodings. For the first position, $\sigma(1)$, we set the causal distribution equal to the non-causal distribution as no extra information is available.
Output. At the causal output, for the prediction of token in position $\sigma(j+1)$, we use a residual connection adding the non-causal hidden state for position $\sigma(j+1)$ e.g. h1 for the token “Speculation” in Figure 1. This ensures the causal target learns to improve over the non-causal distribution. It also aligns the draft and target distributions, increasing the speculative acceptance rate. The causal output from our hybrid architecture depends on both the parameters of the non-causal blocks ($\theta$) and the causal blocks ($\phi$). We write this as $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\bm{x}^{\sigma(i+1:d-1)}))$ specifying that tokens $\bm{x}^{\sigma(1:i)}$ were processed by $\theta$ in the non-causal blocks, whilst future tokens $\phi(\bm{x}^{\sigma(i+1:d-1)})$ were input only into the causal blocks.
Training and Inference. During training, we take a true token sequence, apply masks and input this into the non-causal blocks with the model predicting the true token values in the masked positions. We input the permuted full true token sequence into the causal blocks without masks, with the causal attention operation ensuring each track can only attend to tokens prior to it in the permuted sequence.
During inference, let the currently revealed set of tokens be $\bm{x}^{\sigma(1:i)}$. We pass $\bm{x}^{\sigma(1:i)}$ into the non-causal blocks, with any positions not yet revealed being represented with a mask token. For the causal blocks, which take a full real token sequence as input (no mask tokens), we pass $\bm{x}^{\sigma(1:i)}$ as well as draft tokens generated by the non-causal blocks for future positions, as we will see in Section 3.3.
@@ 3.2 Training Objective
We note that in Equation 8, there are multiple causal distributions that approximate the same true conditional $p_{\rm{data}}(\bm{x}^{\sigma(d)}|\bm{x}^{\sigma(1:d-1)})$ for different values of $i$ which specifies the computation path taken to parameterize $\overset{\rightarrow}{p}_{\theta,\phi}$. We discuss this further in Section 3.3. We train $\theta$ and $\phi$ by jointly maximizing the standard cross entropy loss with a weighting $\frac{D}{D-i}$ that normalizes by the number of masked positions DiD-i (Hoogeboom et al., 2022)
@@
where the expectation is taken over $p_{\rm{data}}(\bm{x})p(\sigma)p(i)$ with $\sigma$ sampled uniformly and $p(i)$ following a noising schedule similar to MDMs with $p(i=D)=0$. The first part of the loss is mathematically equivalent to the MDM loss (Shi et al., 2024; Sahoo et al., 2024), as shown by Ou et al. (2025). The second part of the loss can be seen as a standard AR cross-entropy loss computed over the masked tokens under a random order (Uria et al., 2014). We highlight that evaluating both $\overset{\leftrightarrow}{p}_{\theta}$ and $\overset{\rightarrow}{p}_{\theta,\phi}$ for the objective takes only one single forward pass of the hybrid network. In Appendix E we carry out a FLOP analysis for the extra overhead in our architecture versus a standard transformer, finding only a $0.98\%$ increase in FLOPs for standard settings.
Training curves (Section 5.1) show that the causal distribution $\overset{\rightarrow}{p}_{\theta,\phi}$ achieves a significantly lower loss than the non-causal draft distribution $\smash{\overset{\leftrightarrow}{p}_{\theta}}$, as it can model missing tokens with a non-factorized distribution. We note that our architecture is compatible with both fine-tuning and full-training. In fine-tuning we can freeze the non-causal backbone to a pretrained masked diffusion model and train only the causal blocks on top, as we demonstrate in Section 5.3.